OpenStudy (anonymous):

Find the area of the circle. Use 3.14 for pi. Show your work. Round to the nearest 10th. The diameter is 23yd

OpenStudy (bibby):

Area = pi * r^2 what is r?

OpenStudy (anonymous):

Oops the r is 23

OpenStudy (bibby):

diameter is not the same thing as radius

OpenStudy (anonymous):

I know I ment the radius is 23 not the diameter

OpenStudy (bibby):

ok then plug it into the equation. pi*r*r = pi * 23 *23